Can you please solve this for me? I have to solve this using factoring:
The length of a rectangle is 6cm less than twice its width. Find the dimensions of the rectangle if its area is 108cm2(squared)

Answer by xcentaur(357)

let the width be x
then length=(2x-6)
area is length into breadth
x(2x-6)=108
2x^2-6x-108=0
x^2-3x-54=0
(x-9)(x+6)=0
x=9,-6
width cannot be negative
hence x=9
length=2x-6=18-6=12


xC